# Emergent synchrony in oscillator networks with adaptive arbitrary-order interactions

### v1

@@ Abstract
Dynamics of complex systems are often driven by interactions that extend beyond pairwise links, underscoring the need to establish a correspondence between interpretable system parameters and emergent phenomena in hypergraph-based networks. The current work formulates an adaptive Kuramoto model that incorporates hyperedges of arbitrary order and explores their effects on synchronization. By deriving the exact order parameter dynamics in the thermodynamic limit, analytical expressions governing the collective dynamics are obtained. Subsequent numerics confirm the analytical predictions, in addition to capturing qualitatively different dynamical regimes and phase transitions. Further investigations based on order parameter distributions demonstrate how fluctuations, arising due to finite system size, can influence the long-term system dynamics. These results provide important insights and can have diverse applications, such as designing optimal surgical procedures for drug-resistant epilepsy and identifying the sources of rumours in a social network.
preprint: APS/123-QED
@@
Introduction — Recent studies across different disciplines show that complex systems involve links that extend beyond pairwise edges and include hyperedges that capture interactions between irreducible groups of three or more nodes simultaneously [hypergraphtheoryBretto2013, majhi2022dynamics, shi2022simplicial, boccaletti2023structure, PhysRevLett.131.207401]; see Fig. 1(a) for a sample schematic. Interactions in social dynamics are an archetypal example and include collaborations in scientific research [lungeanu2021team, moore2012analyzing], group interactions in opinion spreading [wang2020social, ghahremani2023novel], and dynamics of consensus [PhysRevE.101.032310, PhysRevE.104.064305], among others. In biological systems, which are often replete with higher-order interactions, neuroscience stands out as a promising domain [giusti2016two, hindriks2025unraveling]. Higher-order connectivity is evident in the human brain [karmelic2022emergent] (also see Fig. 1(b)), and has been linked to avalanches and other neural dynamics [yu2011higher, millan2025spatio, biswas2025effect], and could provide deeper insights into pathological states [PhysRevResearch.5.013074]. These also appear in models describing chemical reactions [wen2023chemical, brauch2013higher], power grids [ghasemi2022higher, PhysRevE.109.024212], and the evolution of ecological systems [swain2022higher, gallardo2024higher, bairey2016high, PhysRevE.111.014309], highlighting their ubiquity across domains.
Figure 1: (a): Schematic representation of hyperedges in the brain: In (i), pink circles indicate nodes, or Broadman areas of functional significance, whereas the lines (edges) and coloured regions (hyperedges) denote pairwise and higher-order interactions, respectively. These can appear at multiple spatial scales, such as macroscopic functional connectivity patterns from fMRI-BOLD time series and microscopic neuronal scale, shown in (ii) and (iii), respectively; (b): Number of interactions in a brain network, as a function of its order, relative to pairwise interactions [karmelic2022emergent, cabral2022metastable]; (c): Temporal variation of intracranial (iEEG) signals from different channels [bougou2025mesoscale, dataset], with the histogram (top-inset) and time series (bottom-inset) of coherence, highlighting its increase during epilepsy, the onset of which is marked with a vertical red line.
The dynamics of systems with higher-order interactions can be described by equations of the form
@@
where $N$ and $D$ denote the number of constituent nodes and the maximum order of interaction, respectively. Oscillations are a characteristic feature often observed in such systems [rabinovich2012oscillations], with their synchronization representing a vital emergent phenomenon having important implications across disciplines [arenas2008synchronization]; for example, epilepsy is often characterized by synchronous activity in the brain [osterhage2007measuring, lehnertz2009synchronization]; see Fig. 1(c) for details. To this end, the Kuramoto model of coupled phase oscillators [RevModPhys.77.137, gupta2014kuramoto] serves as a paradigmatic framework for quantitatively exploring synchronization in a systemic manner. To capture realistic systems, it has also been extensively modified with features such as complex connectivity matrices [jain2023composed, chen2023critical, peron2019onset, tong2018exponential], symmetry-breaking mechanisms [manoranjani2023generalization, biswas2024symmetry], time-delayed interactions [yeung1999time, wu2018dynamics, ross2021dynamics, budzinski2023analytical], higher-order couplings [xu2021spectrum, costa2024bifurcations, costa2025exact, moyal2024rotating, rajwani2023tiered, carballosa2023cluster, sabhahit2024prolonged, skardal2020higher], as well as different combinations of each of these [skardal2022tiered]. These find applications in a myriad of different domains [guo2021overviews, wang2021dynamic, blanter2016kuramoto, fioriti2008stability], and are of particular importance in neuroscience, where they provide a tractable mathematical framework to model complex neuro-oscillatory phenomena [odor2019critical, heggli2019kuramoto, PhysRevLett.120.244101, odor2021effect, pathak2022biophysical, ranjan2024propagation, PhysRevE.111.044310].
@@
This work investigates an adaptive Kuramoto model, which incorporates arbitrary-order interactions of both symmetric and asymmetric types, to explore the effects of higher-order interactions and their interplay with synchronization. The dynamics of the system are assumed to be governed by equations of the form
@@
where $z_{k}=r_{k}e^{{\rm i}\psi_{k}}=N^{-1}\sum_{j=1}^{N}e^{{\rm i}k\phi_{j}}\ \forall\ k\geq 1$ are the Kuramoto-Daido order parameters [clusella2020irregular] and ${\rm i}=\sqrt{-1}$. In this model, the interaction terms are strongly influenced by the corresponding adaptation functions, the inclusion of which represents a fundamental dynamical feature of natural systems [sawicki2023perspectives]. Adaptation allows the characteristic properties of the system to change in response to the collective state of the system and is seen across bio-physical, socio-economic, and engineering systems, and is particularly prevalent in the study of neural activity; for example, spike-frequency adaptation models the reduction in neural firing arising from the fatigue of continuous spiking activity. [PhysRevE.107.024311]. In the context of synchronization, this allows the modulation of the global coupling strength in response to the extent of synchronization in the system [jin2023synchronization], thereby allowing the system to reach a desired state of synchronicity. A universal and well-investigated form of the adaptation function is a power-law, given by $f(r_{1})\propto r_{1}^{\alpha}$, where $\alpha$ controls the nature of the feedback and the resulting dynamics [PhysRevE.75.017201, PhysRevE.102.012219, biswas2024effect].
@@
Analysis — For an infinitely large number of oscillators, an assumption relaxed later, the state of the system can be described by a probability density function $F(\phi,\omega;t)$, which satisfies a continuity equation of the form $\partial_{t}F+\partial_{\phi}(F\dot{\phi})=0$. Since $F(\phi+2\pi,\omega;t)=F(\phi,\omega;t)$, it can be expanded in a Fourier series in $\phi$, where $F_{n}(\omega,t)$ is the $n^{\rm th}$ Fourier coefficient and $F_{-n}(\omega,t)=F^{*}_{n}(\omega,t)$. Using the Ott-Antonsen ansatz [ott2008low, ott2009long], it can be shown that
@@
$\frac{\partial v}{\partial t}+{\rm i}\omega v+\frac{f_{1}(r_{1})\epsilon_{1}}{2}\left(z_{1}e^{-{\rm i}\delta_{1}}v^{2}-z_{1}^{*}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}\delta_{d}}v^{2}-z_{d}^{*}z_{1}^{d-1}e^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}\delta^{\prime}_{d}}v^{2}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}\delta^{\prime}_{d}}\right)\right]=0,$ (3)
where $F_{n}(\omega,t)=v^{n}(\omega,t)$. Subsequently, by assuming that $\omega_{i}$ are sampled from a Cauchy distribution with mean $\omega_{0}$ and scale factor $\Delta$, the order parameters are obtained as $z_{d}=\left[v^{*}(\omega_{0}-i\Delta,t)\right]^{d}$. Substituting into Eq. (3) and re-labelling $z_{1}$ as $z$ for simplicity, it takes the form
@@
$\frac{\partial z}{\partial t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{f_{1}(r)\epsilon_{1}}{2}\left(ze^{-{\rm i}\delta_{1}}-z|z|^{2}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r)\epsilon_{d}}{2}\left(|z|^{2(d-1)}ze^{-{\rm i}\delta_{d}}-|z|^{2d}ze^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r)\epsilon^{\prime}_{d}}{2}\left(z^{2d-1}e^{-{\rm i}\delta^{\prime}_{d}}-|z|^{4}(z^{*})^{2d-3}e^{{\rm i}\delta^{\prime}_{d}}\right)\right].$ (4)
Evidently, the incoherent state is stable only for $\epsilon_{1}\leq\epsilon_{1}^{c}=2\Delta/f_{1}(0)\cos(\delta_{1})$, which denotes the transition point from an asynchronous state to a synchronous state. In contrast, the dynamics of the synchronous states are governed by equations of the form
@@
$\frac{{\rm d}\psi}{{\rm d}t}=S(r,\psi)=\omega_{0}-(1+r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\sin(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\sin(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\sin(\chi^{\prime}_{d})\right)\right],$ (6)
where $\chi^{\prime}_{d}=\delta^{\prime}_{d}-2(d-1)\psi$. These two equations dictate the dynamics of $r(t)$ and $\psi(t)$, the steady state solution $(r_{s},\psi_{s})$, and the subsequent stability analysis provides insights into the overall system dynamics.

### v2

@@ Abstract
Complex physical systems are often governed by interactions that extend beyond pairwise links, underscoring the need to establish a map between interpretable system parameters and emergent synchronisation phenomena in hyper-graphs. To achieve this, the current work formulates an adaptive Kuramoto model that incorporates hyperedges of arbitrary order and explores their effects on synchronisation. By deriving the exact order parameter dynamics in the thermodynamic limit, analytical expressions governing the collective dynamics are obtained. Subsequent numerics confirm the analytical predictions, in addition to capturing qualitatively different dynamical regimes and phase transitions. Further investigations based on numerically constructed order parameter distributions demonstrate how fluctuations due to finite system size can influence the long-term system dynamics by inducing spontaneous transitions. These results provide important insights and can have diverse applications, such as designing optimal surgical procedures for drug-resistant epilepsy in the human brain and identifying the sources of rumours in social networks.
keywords:
Higher-order interactions, adaptive connectivity, explosive phase transitions, synchronization, epilepsy.
journal: Chaos, Solitons & Fractals
[…]
[label1]organization=National Brain Research Centre,addressline=Manesar, city=Gurgaon, postcode=122052, state=Haryana, country=India
1 Introduction
@@
Recent literature across different scientific disciplines shows that real-life complex physical systems involve connections that extend beyond basic pairwise edges and include couplings that capture interactions between groups of three or more nodes simultaneously [hypergraphtheoryBretto2013, majhi2022dynamics, shi2022simplicial, boccaletti2023structure, PhysRevLett.131.207401]; see Fig. 1(a) for a sample schematic. Such interactions, mathematically represented as hyper-edges or simplicial complexes [salnikov2018simplicial], cannot be further decomposed into their constituent pairwise interactions. Systems arising from social dynamics are an archetypal example, and include collaboration networks in scientific research [lungeanu2021team, moore2012analyzing], group interactions in opinion spreading [wang2020social, ghahremani2023novel], and dynamics of consensus [PhysRevE.101.032310, PhysRevE.104.064305], among others. Within biological systems, which are often replete with higher-order interactions, neuroscience stands out as a promising domain [giusti2016two, hindriks2025unraveling]. Higher-order connectivity is evident in the human brain [karmelic2022emergent] (also see Fig. 1(b)), and has been linked to avalanches and other neural dynamics [yu2011higher, millan2025spatio, biswas2025effect], and could provide deeper insights into pathological states [PhysRevResearch.5.013074]. These also appear in models describing chemical reactions [wen2023chemical, brauch2013higher], power grids [ghasemi2022higher, PhysRevE.109.024212], and the evolution of ecological systems [swain2022higher, gallardo2024higher, bairey2016high, PhysRevE.111.014309], highlighting their ubiquity across domains. Mathematically, the dynamics of such systems can be described by
@@
where $N$ and $D$ denote the number of constituent nodes and the maximum order of interaction, respectively. Here, the tensor $A^{(d)}$ and the multivariate function $g^{(d)}(\cdot)$ capture the structure and mechanism of interactions, respectively, of order $d\leq D$.
@@
Oscillations are a characteristic feature often observed in such systems, with their synchronisation representing a vital emergent phenomenon having important implications across disciplines [arenas2008synchronization]. It is particularly important in models of regular and pathological neural dynamics [ansarinasab2025synchronization, ansarinasab2025transition], among which a key example is epilepsy, often characterised by excess synchronous activity between different regions of the brain [osterhage2007measuring, lehnertz2009synchronization]; see Fig. 1(c) for sample schematics. To this end, the Kuramoto model [RevModPhys.77.137, gupta2014kuramoto] serves as a paradigmatic framework for quantitatively exploring synchronisation in a systematic manner. In addition, to capture physically realistic complex systems, it has also been extensively modified to include features such as complex connectivity patterns [jain2023composed, chen2023critical, peron2019onset, tong2018exponential], symmetry-breaking mechanisms [PhysRevE.102.012206, manoranjani2023generalization, biswas2024symmetry], time-delayed interactions [yeung1999time, wu2018dynamics, ross2021dynamics, skardal2022tiered, budzinski2023analytical], phase frustration [buzanello2022matrix, PhysRevE.108.034208, moyal2024rotating, mondal2025enhancing], and higher-order couplings [xu2021spectrum, costa2024bifurcations, costa2025exact, rajwani2023tiered, carballosa2023cluster, sabhahit2024prolonged, skardal2020higher], as well as different combinations of each of these. The Kuramoto model finds applications in a myriad of different domains [guo2021overviews, wang2021dynamic, blanter2016kuramoto, fioriti2008stability], and is of particular importance in neuroscience, where it provides a tractable mathematical framework to capture complex neuro-oscillatory phenomena [odor2019critical, heggli2019kuramoto, PhysRevLett.120.244101, odor2021effect, pathak2022biophysical, ranjan2024propagation, PhysRevE.111.044310]. Additionally, it has also been extended to the so-called “swarmalator” model [o2017oscillators, PhysRevE.111.014209, ansarinasab2024spatial], which introduces additional state variables to capture the dynamics of agents moving through space and interacting through a combination of both phase and spatial interactions, thereby demonstrating a rich repertoire of dynamical states.
Figure 1: (a): Schematic representation of hyperedges in the brain: In (i), pink circles indicate nodes, or Broadman areas of functional significance, whereas the lines (edges) and coloured regions (hyperedges) denote pairwise and higher-order interactions, respectively. These can appear at multiple spatial scales, such as macroscopic functional connectivity patterns from fMRI-BOLD time series and microscopic neuronal scale, shown in (ii) and (iii), respectively; (b): Number of interactions in a brain network, as a function of its order, relative to pairwise interactions; see Refs. [karmelic2022emergent, cabral2022metastable] for more details; (c): Variation of intracranial (iEEG) signals from different channels [bougou2025mesoscale, dataset], with the histogram (top-inset) and time series (bottom-inset) of phase synchrony, highlighting its increase during epilepsy, with its onset marked with a vertical line.
@@
To gain an understanding of synchronisation phenomena, this article investigates an adaptive Kuramoto model that incorporates arbitrary-order interactions of both symmetry-preserving and symmetry-breaking types to explore the effects of higher-order interactions and their interplay with other coupling modalities. The dynamics of the system are assumed to be governed by equations of the form
@@
where $\phi_{i}(t)$ and $\omega_{i}$ represent the instantaneous phase and natural frequency, respectively, of each oscillator $i$. Furthermore,
$z_{k}=r_{k}e^{{\rm i}\psi_{k}}=\frac{1}{N}\sum_{j=1}^{N}e^{{\rm i}k\phi_{j}},\quad k\geq 1,$ (3)
are the complex Kuramoto-Daido order parameters [clusella2020irregular] that quantify the collective dynamics of the systems. In particular, for a nearly incoherent state, $r_{1}$ tends to zero, whereas for a highly synchronous state, $r_{1}$ approaches unity. In this model, the couplings between the individual phases are captured through the second, third, and fourth terms in the RHS of Eq. (2), which represent the pairwise and the two separate types of higher-order interactions, respectively. Evidently, the system is globally coupled, i.e., every node interacts with every other node in an identical manner, across both pairwise and higher-order links.
@@
The interaction terms are strongly influenced by the functions $f_{d}(\cdot)$ and $f_{d}^{\prime}(\cdot)$, the inclusion of which captures adaptivity, a fundamental feature of many natural systems [sawicki2023perspectives], that allows modulating the characteristic properties in response to the current collective state of the overall system. It is seen across bio-physical, socio-economic, and engineering systems, and is particularly prevalent in the study of neural activity; for example, spike-frequency adaptation models the reduction in neural firing arising from the fatigue of continuous spiking activity [PhysRevE.107.024311]. In the context of synchronisation, this allows the modulation of the global coupling strengths (i.e., $\epsilon_{d}$ and $\epsilon_{d}^{\prime}$) in response to the extent of synchronisation in the system [jin2023synchronization], quantified through $r_{1}$, thereby allowing the system to reach a desired state of synchronicity. For example, a universal and well-investigated form of the adaptation function is a power-law, where the exponent controls the nature of the feedback and the resulting dynamics [PhysRevE.75.017201, PhysRevE.102.012219, biswas2024effect], a variation of which is investigated in later sections. Furthermore, the coupling terms also comprise of phase lags in each order of interaction (i.e., $\delta_{d}$ and $\delta_{d}^{\prime}$), the primary effect of which is to modify the common frequency to some desired value [lohe2015synchronization], whereas both symmetry-preserving and symmetry-breaking higher-order interactions terms are introduced to capture non-trivial generalizations of the basic Kuramoto model, such as the Winfree model [manoranjani2023generalization, PhysRevE.96.042208].
The rest of the article is arranged as follows: Sec. 2 outlines the dimensionality reduction of the proposed system, which provides analytical insights into the overall dynamics. In Sec. 3, these results are validated against numerical simulations in Sec. 3.1, highlighting different dynamical regimes, with subsequent subsections showcasing the relative effect of interactions (Sec. 3.2) and the effect of fluctuations in finite-sized networks (Sec. 3.3). The paper ends with a summary of the key results, limitations, and future research directions in Sec. 4.
2 Analytical Insights
To obtain analytical insights into the dynamics of the proposed system, Eq. (2) is first rewritten, using the definition of the order parameter $z_{k}$ in Eq. (3), in the form
$\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r_{1})\epsilon_{1}}{2{\rm i}}\left(z_{1}e^{-{\rm i}(\delta_{1}+\phi_{i})}-z_{1}^{*}e^{{\rm i}(\delta_{1}+\phi_{i})}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2{\rm i}}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}(\delta_{d}+\phi_{i})}-z^{*}_{d}z_{1}^{d-1}e^{{\rm i}(\delta_{d}+\phi_{i})}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2{\rm i}}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}(\delta^{\prime}_{d}+\phi_{i})}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}(\delta^{\prime}_{d}+\phi_{i})}\right)\right],$ (4)
@@ 2 Analytical Insights
where * represents complex conjugation. For an infinitely large number of oscillators, an assumption which is relaxed in the later sections, the state of the system in Eq. (2) can be described by a joint probability density function $F(\phi,\omega;t)$, which satisfies a continuity equation of the form
$\frac{\partial F}{\partial t}+\frac{\partial(F\dot{\phi})}{\partial\phi}=0$ (5)
and the condition
$\int_{0}^{2\pi}F(\phi,\omega;t)\ {\rm d}\phi=g(\omega),$ (6)
where $\dot{\phi}$ is given by Eq. (4). Since $F(\phi+2\pi,\omega;t)=F(\phi,\omega;t)$, it can also be expanded as a series of the form
$F(\phi,\omega;t)=\frac{g(\omega)}{2\pi}\left[1+\sum_{n=1}^{\infty}\left(F_{n}e^{{\rm i}n\phi}+F^{*}_{n}e^{-{\rm i}n\phi}\right)\right],$ (7)
where $F_{n}\equiv F_{n}(\omega,t)$ is the $n^{\rm th}$ Fourier coefficient. By using the Ott-Antonsen ansatz [ott2008low, ott2009long], the Fourier modes are assumed to be of the form $F_{n}(\omega,t)=v^{n}(\omega,t)$, with the restriction $|v(\omega,t)|<1$ so that the sum in Eq. (7) converges. This implies that Eq. (5) can be re-written in terms of $v(\omega,t)$ as
@@
$\frac{\partial v}{\partial t}+i\omega v+\frac{f_{1}(r_{1})\epsilon_{1}}{2}\left(z_{1}e^{-{\rm i}\delta_{1}}v^{2}-z_{1}^{*}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}\delta_{d}}v^{2}-z_{d}^{*}z_{1}^{d-1}e^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}\delta^{\prime}_{d}}v^{2}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}\delta^{\prime}_{d}}\right)\right]=0.$ (8)
Assuming that the natural frequencies $\omega_{i}$ are sampled from a Cauchy distribution with mean $\omega_{0}$ and scale factor $\Delta$, the summation in Eq. (3) can be replaced with an integral, which can then be evaluated to obtain
$z_{d}=\int_{0}^{2\pi}\int_{-\infty}^{\infty}e^{{\rm i}d\phi}F(\phi,\omega;t)\ {\rm d}\omega\ {\rm d}\phi=\Big[v^{*}(\omega_{0}-i\Delta,t)\Big]^{d}.$ (9)
Upon by substituting Eq. (9) into Eq. (8) and re-labelling $z_{1}$ as $z$ for the sake of simpler notation, Eq. (8) takes the form
$\frac{{\rm d}z}{{\rm d}t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{f_{1}(r)\epsilon_{1}}{2}\left(ze^{-{\rm i}\delta_{1}}-z|z|^{2}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r)\epsilon_{d}}{2}\left(|z|^{2(d-1)}ze^{-{\rm i}\delta_{d}}-|z|^{2d}ze^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r)\epsilon^{\prime}_{d}}{2}\left(z^{2d-1}e^{-{\rm i}\delta^{\prime}_{d}}-|z|^{4}(z^{*})^{2d-3}e^{{\rm i}\delta^{\prime}_{d}}\right)\right],$ (10)
yielding a two-dimensional coupled system, the analysis of which is simpler compared to Eq. (2). First, the stability of the incoherent state of the system can be obtained by linearizing Eq. (10) about $z(t)=0$, leading to the equation
$\frac{{\rm d}z}{{\rm d}t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{z}{2}f_{1}(0)\epsilon_{1}e^{-{\rm i}\delta_{1}}.$ (11)
This implies that $z=0$ is stable only for
$\epsilon_{1}\leq\epsilon_{1}^{c}=\frac{2\Delta}{f_{1}(0)\cos(\delta_{1})},$ (12)
which corresponds to the critical point beyond which the system undergoes a phase transition from an asynchronous state to a synchronous state. Alternatively, Eq. (10) can also be rewritten as
@@
$\begin{split}\frac{{\rm d}r}{{\rm d}t}+{\rm i}r\frac{{\rm d}\psi}{{\rm d}t}&={\rm i}(\omega_{0}+{\rm i}\Delta)r+\frac{rf_{1}(r)\epsilon_{1}}{2}H(r,\delta_{1})+\\ &\sum_{d=2}^{D}r^{2d-1}\left[\frac{f_{d}(r)\epsilon_{d}}{2}H(r,\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}H(r,\chi^{\prime}_{d})\right],\end{split}$ (13)
where $H(x,y)=(1-x^{2})\cos(y)-{\rm i}(1+x^{2})\sin(y)$ and $\chi^{\prime}_{d}=\delta^{\prime}_{d}-2(d-1)\psi$. Therefore, the dynamics of the synchronous states (i.e., where $r>0$) are governed by coupled differential equations of the form
@@
$\frac{{\rm d}\psi}{{\rm d}t}=\Psi(r,\psi)=\omega_{0}-(1+r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\sin(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\sin(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\sin(\chi^{\prime}_{d})\right)\right],$ (15)
which can be obtained by separating the real and imaginary parts of Eq. (13). The steady state solutions of the above two equations (say, $(r_{s},\psi_{s})$) provide insights into the system dynamics. These can be obtained as solutions of the algebraic equations $R(r_{s},\psi_{s})=0$ and $\Psi(r_{s},\psi_{s})=0$, which can be expressed as
$\left(\frac{\Delta}{1-r^{2}_{s}}\right)-\sum_{n=1}^{D}\sigma_{n}(r_{s})\cos(\delta_{n})=\sum_{m=2}^{D}\sigma_{m}^{\prime}(r_{s})\cos(\chi_{m,s}^{\prime}),$ (16)
and
$\left(\frac{\omega_{0}}{1+r^{2}_{s}}\right)-\sum_{n=1}^{D}\sigma_{n}(r_{s})\sin(\delta_{n})=\sum_{m=2}^{D}\sigma_{m}^{\prime}(r_{s})\sin(\chi_{m,s}^{\prime}),$ (17)
where $\sigma_{n}(r)=f_{n}(r)\epsilon_{n}r^{2(n-1)}/2$, $\sigma^{\prime}_{n}(r)=f^{\prime}_{n}(r)\epsilon^{\prime}_{n}r^{2(n-1)}/2$, and $\chi^{\prime}_{m,s}$ denotes $\chi^{\prime}_{m}$ evaluated at $\psi={\psi_{s}}$. Obtaining their stability entails calculating the Jacobian $J$ of Eqs. (14) and (15), the analytical expressions of its elements $J_{ij}$ are listed in Eqs. (22)-(25) of A.
Therefore, taken together, the analytical results presented in this section provide a theory of synchronization in systems with arbitrary-order interactions, while also accounting for the effects of different adaptive mechanisms, phase lags, and generalized forms of interactions that break rotational symmetry in the Kuramoto model. The subsequent sections are devoted to verifying these results through numerical simulations, as well as investigating other properties of the proposed system through numerical simulations.
3 Numerical Results
Although the analytical results derived in the previous section constitute a general, arbitrary-order theoretical framework, further analytical progress is feasible only under specific simplifying assumptions on the parameter values, whereas incorporating arbitrarily large orders of interactions would increase simulation times beyond reasonable limits. Therefore, to validate the obtained analytical expressions, the simplest yet representative example of Eq. (2) is chosen. It corresponds to a system consisting of pairwise and triadic interactions, similar to that reported in Ref. [biswas2024symmetry], and is described by equations of the form
$\begin{split}\frac{{\rm d}\phi_{i}}{{\rm d}t}=&\omega_{i}+\frac{f_{1}(r)\epsilon_{1}}{N}\sum_{j=1}^{N}\sin(\phi_{j}-\phi_{i})+\\ &\frac{f_{2}(r)\epsilon_{2}}{N^{2}}\sum_{j,k=1}^{N}\Bigg[\sin(2\phi_{k}-\phi_{j}-\phi_{i})+\xi\sin(2\phi_{k}+\phi_{j}-\phi_{i})\Bigg],\end{split}$ (18)
where $\xi=\epsilon_{2}^{\prime}/\epsilon_{2}$. For the sake of simplicity, $f_{d}(r)$ are assumed to be of the form $f_{d}(r)=(1+\gamma r)^{\pm 2}$ for $d=1$ and $d=2$, respectively, whereas $\gamma$ is a constant that generates a family of qualitatively different functional forms; also see Fig. 2.
(a) $d=1$
(b) $d=2$
Figure 2: Plot of $f_{d}(r)$ as a function of $r\in[0,1]$ for (a) $d=1$ and (b) $d=2$, where the colours denote different values of $\gamma$.
This choice represents the simplest possible functional form of a polynomial adaptation function, which has previously been shown to induce both continuous and explosive phase transitions [biswas2024effect], and has been set up such that it introduces qualitatively different adaptation behaviour (i.e., increasing and decreasing) across the two orders of interactions.